# supplementary

### v1

@@ 3 Discussion
Generative AI voting: an inevitable race to safeguard democracy. Generative AI voting is likely to emerge as an inevitable technological convergence of AI and e-voting solutions currently adopted in real-world. On the one hand, generative AI and large language models are expected to become more open, pervasive and accessible to citizens [45, 15]. AI personal assistants are already part of everyday life [46, 47, 48], with their generative version expected to follow. On the other hand, the mandate of more direct, secure and active participation in decision making for public matters is expected to further scale up electronic voting solutions and digital platforms. For instance, participatory budgeting elections are mainly conducted digitally, while Estonia has already institutionalized a digital identity for 99% of its citizens as well as electronic voting since 2005 [49]. As Toomas H. Ilves, former president of Estonia, emphasized "with the digital signature and the machine-readable ID card, we created the e-citizen". In the light of these converging advancements, the findings of this study become spot on to understand the implications of such future, while they are significant to prepare safeguards for digital democracy.
@@
Fair voting design as a safeguard for generative AI voting. Large language models prove not ready yet to accurately represent humans in complex voting scenarios such as participatory budgeting. However, a striking finding of this study is that consistency of large language models becomes significant under fair voting aggregation methods that promote a proportional representation of voters’ preferences. This motivates a huge opportunity to get democracy "right" in the digital era of AI: move to alternative voting methods that yield fairer voting outcomes for all, while shielding democratic outcomes from AI biases and inconsistencies. The momentum for this win-win paradigm shift is tremendous as fair voting rules such as the method of equal shares have been recently tested in real-world with grand success: more winning projects with higher proportional representation of citizens’ preferences [39, 22]. How to scale up these democratic blueprints remains an open question, but AI can play a key role here.
@@
Why fair collective choice is resilient to AI biases and inconsistencies. We provide a plausible explanation of this significant and striking finding. There is evidence that the method of equal shares has an inherent stability in the resulting voting outcomes [22]. For instance, even with 25-50% of the voters, the elected projects do not significantly change [22]. This is despite the fact that projects may end with very different number of total votes, as low and middle-cost projects require a very minimal support to get elected, and as a result, these winning projects are likely to be retained in the winning set, even with different choices or set of voters. Such projects are expected to be an origin of consistency. Indeed, this effect is also observed in the real-world voting scenario of City Idea, as with 80% abstaining voters, 84% of the winners are retained with equal shares, see Figure LABEL:fig:mixed_populations-abstaining. As City Idea promoted equal shares already in the project ideation phase and made apriori use of equal shares for the aggregation of the ballots, this study comes with significant realism and merit for the validity of the findings. This comes in stark contrast to other earlier studies [22, 39] that hypothesize the application of equal shares over proposed projects and ballots aggregated with the standard method of utilitarian greedy.
@@
Trustworthy generative AI voting: a call for research and policy action. What information large language models use to reason about voting decisions is influential for different types biases to manifest. Unconscious and conformity biases, related to political belief, race and environment, reinforce human choices or yield inconsistencies. Time discounting biases, related to the support of projects on welfare, public space and sport, explain preference transitivity among different ballot formats. These significant findings demonstrate that training data in generative AI voting are expected to be highly impactful for the representation of the voting outcomes in the population. Ethical and democratic guidelines are urgently needed, particularly for the use of (generative) AI in voting processes. For instance, who shall determine the input training data of AI representatives? Should the training data involve only self-determined personal information of voters, or shall these be augmented with more universal knowledge and experts’ opinions? How to protect the privacy and autonomy of voters when training such AI representatives [48]? Will citizens retain power to control AI representatives that reflect their values and beliefs, while remaining accountable? These are some key questions as a basis of a call for action on research and policy in an emerging era of generative AI voting.
@@ 4 Methods
We show here how AI representatives are emulated and the real-world data based on which several voting scenarios are constructed. We also illustrate the evaluation approach in detail and the studied human cognitive biases. Finally, the approach to explain the inconsistencies and biases of generative AI is outlined.
@@ 4.1 Emulating AI representatives
    Ballot formats: The two participatory budgeting scenarios are either approvals (n-approval ballot where ‘n’ is the number of projects approved) or cardinal preferences (score or cumulative) [50, 51]. In the case of score ballots, the voters provide a score to a project from a specified range such as 1 to 5 with 1 being the lowest preference and 5 being the highest. In the case of cumulative ballots, the voters distribute a number of given points among the projects. Single-choice ballots are derived from single approvals, top scores or top points distributed in a project.
@@
Generative AI. The three inputs determine a three-tiered context for ballot generation. This provides comprehensive information about the voter and the voting process to prompt large language models for choice generation without any additional fine-tuning process. We use proprietary models such as GPT 3.5, GPT3 [52] and the open-source Llama2-13B [53]. GPT 3.5 and GPT 4 have similar performance for these tasks [54], hence we resort to GPT 3.5 for the analysis. Furthermore, both GPT 3.5 and GPT 4 use the “Let’s think step by step” to evaluate a prompt and find a solution that mimics human reasoning [54]. For each of the large language model, predictions are made based on the five combinations of personal human traits (see Figure 1), however the predictions with all models have the consistency to human choices, which we use for further analysis.
@@
Predictive AI benchmark. The personal human traits are used as features to predict the ballots using classification methods [55]. We sample 40% of the data and use synthetic minority oversampling [56] to augment the data. This 40% is decided empirically based on the performance of the neural network models. We also divide the prediction into multiple sub-problems, which predict the approvals or assigned scores for every option and then compile the final ballot. This model is then used to predict the 60% of the unknown ballots and the 40% of the known ballots. As such, we were able to train a model with good precision and recall using 40% of the data and sampling, which could be attributed to clusters of voters generating ballots of similar patterns.
@@
Ballot aggregation. For the US elections, majority determines the winner. For the two voting scenarios of participatory budgeting, the generated ballots by the AI representatives are aggregated using either the utilitarian greedy method or the method of equal shares. Utilitarian greedy simply selects the next most popular project, the one with the most received votes, as long as the available budget is not exhausted. Equal shares provides a proportional representation of voters’ preferences by dividing the budget equally among voters as endowments. This part of the budget can only be used to fund projects that the voter has voted for. The method goes through all project options, beginning with the projects with the highest number of votes. It selects a project if it can be funded using the budget shares of those who voted for the project. A full explanation of the is out of the scope of this article and can be found in earlier work [38, 50, 51]. In practice, the method of equal shares is likely to sacrifice an expensive popular project to select instead several low-cost projects, which, in combination, satisfy the preferences of more voters [39, 22]. As a result, equal shares tends to elect more winning projects. Because of this effect, it likely that consistency measurements based on the Jaccard similarity yield higher values for equal shares. This is the reason we control for the number of winning projects in equal shares by counting a subset of the most popular winning projects which is equal in number with winners of greedy. .
@@
Emulated voting processes with human voters and AI representatives. We emulate the following voting processes: (i) Human voters exclusively with 100% turnout. (ii) human voters exclusively with varying turnout in the range [10%,100%] with a step of 10%. (iii) Mixed populations of human voters and AI representatives assessed whether they can recover the consistency of voting outcomes lost by different human voters turnouts in the range [10%,100%] with a step of 10%. The ratio of AI representatives for the abstained voters varies in the range [25%,100%] with a step of 25%. (iv) AI representatives exclusively.
@@
Data collection infrastructure. Generative AI choices are collected via API prompts to large language models from 16th of June 2023 to 8th of November 2023. We prompted the large language models using the zero-shot learning feature [52], which does not require any specific fine-tuning. We use chain of thought prompting [57] to provide a comprehensive and systematic flow of information for better interpretability. The large language models are run with temperature settings of 0.2 to 0 (20 runs for each temperature setting), for a focused decision with less randomness [58]. As we are dealing with a significantly large decision space particularly the 33 projects in the actual voting, running with a high-temperature setting can lead to random decision-making by the large language models. To maintain consistency, we retain the projects that are approved at the deterministic temperature setting (the responses are easily reproducible at temperature setting = 0). Further, if a project is selected both in 0.1 and 0.2, we include the approval for that project in the ballot to also incorporate diversity. In the case of scores, we take the frequency of the distinct scores assigned to a project across different temperature setting and add the score with the highest frequency. However, for the actual City Idea dataset for which we emulate cumulative voting, where the total score provided in a ballot should be less than a maximum point. Hence it would not be not be accurate to derive a score for a project from multiple outputs corresponding to different temperature settings, hence we conducted experiments only with temperature setting = 0.2. The decoding algorithm for the large language model is set to the nucleus sampling approach [57]. Compared to purely random sampling, it avoids generating outliers or improbable sequences as choices.
For the predictive AI choices, we use neural networks with 3 hidden layers and relu [59] activation function for the sub-networks in the neural network. We employ the ensemble method [60] of designing multiple neural networks with different parameters and merging in the final layer with cross entropy loss [61]. This helps to generalize on different sizes of datasets and improve the prediction scores on unseen data.
@@ 4.2 Voting datasets
US elections. The 2012, 2016, and 2020 survey waves of the American National Election Study (ANES) [62] are used. They encompass 4,500 votes together with the respective voter information for each of these three years. This mainly includes socio-demographic characteristics: (i) racial/ethnic self-identification, (ii) gender, (iii) age, (iv) conservative-liberal ideological self-placement, (v) party identification (political belief), (vi) political interest, (vii) church attendance, (viii) discussion of politics with family/friends, (ix) feelings of patriotism associated with the American flag, and (x) state of residence. A total of 6 elections are emulated using Llama-13B and GPT-3.5 based on single-choice ballots, with winners decided by majority for the three years. We extracted information from 20,650 voters of the ANES dataset [62] and after pre-processing (removing voters with missing values in more than 5 of the personal traits), a set of 17,010 voters is used. The examples of the prompts used to generate the AI choices can be found in Table LABEL:table:Anes.
City Idea: Participatory budgeting campaign in Aarau, Switzerland. The data from a recent innovative participatory budgeting field study are used [32, 39], which was conducted with ethical approval from University of Fribourg (#2021-680). It run 2023 and is rigorously designed to assess the application of equal shares for the first time in real world, in combination with cumulative voting, using the open-source Stanford Participatory Budgeting platform [33]. As such, the field study includes a survey conducted before voting linking the choices of survey respondents and voters. We use the following personal human traits from the survey (more information in Tables LABEL:table:prologue1 and LABEL:table:prologue2): (i) 9 fundamental socio-demographic characteristics (e.g., age, citizenship, education) and 2 political interests (political beliefs and trust in democracy), (ii) preferences for 9 different types of projects and 6 beneficiaries and (iii) 4 types of preferences for qualities of the voting outcome.
Table 1: The studied dimensions across three real-world voting scenarios. They provide the necessary diversity to generalize the findings of this study as they include different scales in number of voters, different ballot formats and aggregation methods, low and high numbers of alternatives, different personal human traits for studying a broad spectrum of biases and both generative and predictive AI methods.
@@
City Idea includes a mock voting of hypothetical costed projects as part of the initial survey before the actual voting took place. The AI representatives are constructed using the same personal human traits to emulate both mock and actual voting. More information about the two voting scenarios is illustrated below:
@@
    Mock Voting: Five hypothetical costed projects belonging in different categories are put for mock voting as part of the initial survey. Table LABEL:table:prologueS illustrates the project alternatives and their cost. The choice of 3134 voters over the same alternatives is tested with three different ballot formats in a sequence, starting with the simplest one of single choice to the most complex ones of approvals and score voting. This allows us to emulate a total of 45 elections based on various combinations of the ballot format and personal human traits using the AI models and the aggregation methods outlined in Figure 1. We also emulate an election based on the predictive AI model.
@@
    Actual Voting: Using the Stanford Participatory Budgeting platform [33], 1703 voters cast their vote using cumulative ballots by distributing 10 points to at least 3 projects of their preference, out of 33 projects in total (see Table LABEL:table:prologueA for project descriptions). A subset of 505 of these voters, which participated in the initial survey and provided their personal human traits, are used to construct the AI representatives. The ballot formats of single choice and approvals are derived from the cumulative ballots by taking the project with the most points and the projects that received any point respectively. This allows us to emulate a total of 27 elections based on various combinations of the ballot formats and personal human traits using the AI models and the aggregation methods outlined in Figure 1. We have also emulated an election based on the predictive AI model.
@@ 4.3 Evaluation of choices by AI representatives
We compare the emulated elections with AI representatives with the real-world elections of human voters at two levels: (i) individual choice, i.e. the ballots, and (ii) collective choice, i.e. the resulting voting outcomes. Consistency is the main measure of assessment.
@@ 4.4 Human cognitive biases in AI collective decision making
Surrogation biases. This reflects how humans favor simpler measures to assess the impact over ones that are more precise and harder to evaluate. Korteling et al. [41] argue these biases manifest when deciding projects with large societal impact such as health or education, while their outcome is subject of different satisfaction levels among citizens. The project outcomes may be perceived as successful by part of society using easy-to-evaluate metrics instead of looking at long-term effects on the community. For instance, a timely vaccination drive may be preferred over significant changes to vaccination protocols or health insurance policies covering vaccination. Hence these projects are likely to be preferred more as they are likely to come with more intuitive ways to assess for the broader population. This is reflected with average winning rate of the health and education-related projects of 38.1% and 36.2% respectively in Poland [39], where participatory campaigns have been actively hosted in the last decade.

### v2

@@ 3 Discussion
Generative AI voting: an inevitable race to safeguard democracy. Generative AI voting is likely to emerge as an inevitable technological convergence of AI and e-voting solutions currently adopted in real-world. On the one hand, generative AI and large language models are expected to become more open, pervasive and accessible to citizens [43, 15]. AI personal assistants are already part of everyday life [44, 45, 46], with their generative version expected to follow. On the other hand, the mandate of more direct, secure and active participation in decision making for public matters is expected to further scale up electronic voting solutions and digital platforms. For instance, participatory budgeting elections are mainly conducted digitally, while Estonia has already institutionalized a digital identity for 99% of its citizens as well as electronic voting since 2005 [47]. As Toomas H. Ilves, former president of Estonia, emphasized "with the digital signature and the machine-readable ID card, we created the e-citizen". In the light of these converging advancements, the inter-operation of a generative personal voting assistant with a digital voting platform becomes technologically feasible along with the citizens’ need to have a more direct saying in several public matters and consultations. Therefore, the findings of this study become spot on to understand the implications of such a future, while they are significant to prepare safeguards for digital democracy.
@@
Fair voting design as a safeguard for generative AI voting. Large language models prove not ready yet to accurately represent humans in complex voting scenarios such as participatory budgeting. However, a striking finding of this study, built upon our demonstrated real-world evidence in Aarau, is that consistency of large language models becomes significant under fair voting aggregation methods that promote a proportional representation of voters’ preferences. This motivates a huge opportunity to get democracy "right" in the digital era of AI: move to alternative voting methods that yield fairer voting outcomes for all, while shielding democratic outcomes from AI biases and inconsistencies. The momentum for this win-win paradigm shift is tremendous, which would not be possible otherwise to demonstrate without the real-world empirical evidence based on which this study unfolds: fair voting rules such as the method of equal shares promotes under-represented projects and yield more winning projects with higher proportional representation of citizens’ preferences [39, 22]. Nevertheless, how to scale up these democratic blueprints remains an open question, but AI can play a key role here.
@@
Why fair collective choice is resilient to AI biases and inconsistencies. We provide an explanation of this significant and striking finding. There is evidence that the method of equal shares has an inherent stability in the resulting voting outcomes [22]. For instance, even with 25-50% of the voters, the elected projects do not significantly change [22]. This is despite the fact that projects may end with very different number of total votes, as low and middle-cost projects require a very minimal support to get elected, and as a result, these winning projects are likely to be retained in the winning set, even with different choices or set of voters. Such projects are expected to be an origin of consistency. Indeed, this effect is also observed in the real-world voting scenario of City Idea, as with 80% abstaining voters, 84% of the winners are retained with equal shares, see Figure LABEL:fig:mixed_populations-abstaining. As City Idea promoted equal shares already in the project ideation phase and made apriori use of equal shares for the aggregation of the ballots, this study becomes the first of its kind and comes with compelling realism and merit for the validity of our significant findings. This comes in stark contrast to other earlier studies [22, 39] that hypothesize the application of equal shares over proposed projects and ballots aggregated with the standard method of utilitarian greedy.
@@
Trustworthy generative AI voting: a call for research and policy action. What information large language models use to reason about voting decisions is influential for different types biases to manifest. Unconscious and conformity biases, related to political belief, race and environment reinforce human choices or yield inconsistencies. Time discounting biases, related to the support of projects on welfare and sport, explain preference transitivity among different ballot formats. These significant findings demonstrate that training data in generative AI voting are expected to be highly impactful for the representation of the voting outcomes in the population. Ethical and democratic guidelines are urgently needed, particularly for the use of (generative) AI in voting processes. For instance, who shall determine the input training data of AI representatives? Should the training data involve only self-determined personal information of voters, or shall these be augmented with more universal knowledge and experts’ opinions? How to protect the privacy and autonomy of voters when training such AI representatives [46]? Will citizens retain power to control AI representatives that reflect their values and beliefs, while remaining accountable? These are some key questions as a basis of a call for action on research and policy in an emerging era of generative AI voting.
@@ 4 Methods
We show here how AI representatives are emulated and the real-world data based on which the voting scenarios are constructed. We also illustrate the evaluation approach and the studied human cognitive biases. Finally, the approach to explain the inconsistencies and biases of generative AI is outlined.
@@ 4.1 Emulating AI representatives
    Ballot formats: The two participatory budgeting scenarios are either approvals (n-approval ballot where ‘n’ is the number of projects approved) or cardinal preferences (score or cumulative) [48, 49]. In the case of score ballots, the voters provide a score to a project from a specified range such as 1 to 5 with 1 being the lowest preference and 5 being the highest. In the case of cumulative ballots, the voters distribute a number of given points among the projects. Single-choice ballots are derived from single approvals, top scores or top points distributed to a project.
@@
Generative AI. We use proprietary models such as GPT 3.5, GPT3 [50] and the open-source Llama2-13B [51]. GPT 3.5 and GPT 4 have similar performance for these tasks [52], hence we resort to GPT 3.5 for the analysis. Furthermore, both GPT 3.5 and GPT 4 use the “Let’s think step by step” to evaluate a prompt and find a solution that mimics human reasoning [52]. For each of the large language models, predictions are made based on the five combinations of personal human traits (see Figure 1), however the predictions with all models have the highest consistency to human choices, which we use for the analysis.
@@
Predictive AI benchmark. The personal human traits are set as features to predict the ballots using classification methods [53]. We sample 40% of the data and use synthetic minority oversampling [54] to augment the data. This 40% is decided empirically based on the performance of the neural network models. We also divide the prediction into multiple sub-problems, which predict the approvals or assigned scores for every option and then compile the final ballot. This model is then used to predict 60% of the unknown ballots and 40% of the known ballots.
@@
Ballot aggregation. For the US elections, majority determines the winner. For the two voting scenarios of participatory budgeting, the generated ballots by the AI representatives are aggregated using either the utilitarian greedy method or the method of equal shares. Utilitarian greedy simply selects the next most popular project, the one with the most received votes, as long as the available budget is not exhausted. Equal shares provides a proportional representation of voters’ preferences by dividing the budget equally among voters as endowments. This part of the budget can only be used to fund projects that the voter has voted for. The method goes through all project options, beginning from the projects with the highest number of votes. It selects a project if it can be funded using the budget shares of those who voted for the project. A full explanation of the is out of the scope of this article and can be found in earlier work [38, 48, 49]. In practice, the method of equal shares is likely to sacrifice an expensive popular project to select instead several low-cost projects, which, in combination, satisfy the preferences of more voters [39, 22]. As a result, equal shares tends to elect more winning projects. Because of this effect, it likely that consistency measurements based on the Jaccard similarity yield higher values for equal shares. This is the reason we control for the number of winning projects in equal shares by counting a subset of the most popular winning projects which is equal in number with the winners of the utilitarian greedy method.
@@
Emulated voting processes with human voters and AI representatives. We emulate the following voting processes: (i) Human voters exclusively with 100% turnout. (ii) Human voters exclusively with varying turnout in the range [10%,100%] with a step of 10%. (iii) Mixed populations of human voters and AI representatives assessed of whether they can recover the consistency of voting outcomes lost by different human voters turnouts in the range [10%,100%] with a step of 10%. The ratio of AI representatives for the abstained voters varies in the range [25%,100%] with a step of 25%. (iv) AI representatives exclusively.
@@
Data collection infrastructure. Generative AI choices are collected via API prompts to large language models from 16th of June 2023 to 8th of November 2023. We prompted the large language models using the zero-shot learning feature [50], which does not require any specific fine-tuning. We use chain of thought prompting [55] to provide a comprehensive and systematic flow of information for better interpretability. The large language models run with temperature settings of 0.2 to 0 (20 runs for each temperature setting), for a focused decision with less randomness [56]. As we are dealing with a significantly large decision space, particularly the 33 projects in the actual voting, running with a high-temperature setting can lead to random decision-making. To maintain consistency, we retain the projects that are approved at the deterministic temperature setting (the responses are easily reproducible at temperature setting = 0). Furthermore, if a project is selected both in 0.1 and 0.2, we include the approval for that project in the ballot to also incorporate diversity. In the case of score voting, we take the frequency of the distinct scores assigned to a project across different temperature setting and add the score with the highest frequency. However, for the actual City Idea scenario for which we emulate cumulative voting, the total score provided in a ballot should be lower than the total number of points. Therefore, it would not be be accurate to derive a score for a project from multiple outputs corresponding to different temperature settings, hence we conducted experiments only with the temperature setting of 0.2. The decoding algorithm for the large language model is set to the nucleus sampling approach [55]. Compared to purely random sampling, it avoids generating outliers or improbable sequences of choices.
For the predictive AI choices, we use neural networks with 3 hidden layers and the relu [57] activation function for the sub-networks in the neural network. We employ the ensemble method [58] of designing multiple neural networks with different parameters and merging in the final layer with cross entropy loss [59]. This helps to generalize on different sizes of datasets and improve the prediction scores on unseen data.
@@ 4.2 Voting datasets
US elections. The 2012, 2016, and 2020 survey waves of the American National Election Study (ANES) [60] are used. They encompass 4,500 votes together with the respective voter information for each of these three years. This mainly includes socio-demographic characteristics: (i) racial/ethnic self-identification, (ii) gender, (iii) age, (iv) conservative-liberal ideological self-placement, (v) party identification (political belief), (vi) political interest, (vii) church attendance, (viii) discussion of politics with family/friends, (ix) feelings of patriotism associated with the American flag, and (x) state of residence. A total of 9 elections are emulated using Llama, GPT-3.5 and the predictive ML model based on single-choice ballots, with winners decided by majority for the three years. We extracted information from 20,650 voters of the ANES dataset [60] and after pre-processing (removing voters with missing values in more than 5 of the personal traits), a set of 17,010 voters is used. The examples of the prompts used to generate the AI choices can be found in Table LABEL:table:Anes.
City Idea: Participatory budgeting campaign in Aarau, Switzerland. The data from a recent innovative participatory budgeting field study are used [32, 39], which was conducted with ethical approval from University of Fribourg (#2021-680). It run in 2023 and is rigorously designed to assess the application of equal shares for the first time in real world, in combination with cumulative voting, using the open-source Stanford Participatory Budgeting platform [33]. As such, the field study includes a survey conducted before voting linking the choices of survey respondents and voters. We use the following personal human traits from the survey (more information in Tables LABEL:table:prologue1 and LABEL:table:prologue2): (i) 9 key socio-demographic characteristics (e.g., age, citizenship, education) and 2 political interests (political beliefs and trust in democracy), (ii) preferences for 9 different types of projects and 6 beneficiaries and (iii) 4 types of preferences for qualities of the voting outcome.
Table 1: The studied dimensions across three real-world voting scenarios. They provide the necessary diversity to generalize the findings of this study as they include different scales in number of voters, different ballot formats and aggregation methods, low and high numbers of alternatives, different personal human traits for studying a broad spectrum of biases including both generative and predictive AI methods.
@@
Emulated elections 9 91 27
Two participatory budgeting voting scenarios are studied in the context of the real-world campaign of City Idea:
@@
    Mock Voting: Five hypothetical costed projects belonging in different categories are put for mock voting as part of the initial survey. Table LABEL:table:prologueS illustrates the project alternatives and their cost. The choice of 3134 voters over the same alternatives is tested with three different ballot formats in a sequence, starting with the simplest one of single choice to the most complex ones of approvals and score voting. This allows us to emulate a total of 45 elections = 3 ballot formats x 2 ballot aggregation methods x 3 AI models x 5 personal traits + 1 for the predictive ML model.
@@
    Actual Voting: Using the Stanford Participatory Budgeting platform [33], 1703 voters cast their vote using cumulative ballots by distributing 10 points to at least 3 projects of their preference, out of 33 projects in total (see Table LABEL:table:prologueA for project descriptions). A subset of 505 of these voters, which participated in the initial survey and provided their personal human traits, are used to construct the AI representatives. The ballot formats of single choice and approvals are derived from the cumulative ballots by taking the project with the most points and the projects that received any point respectively. This allows us to emulate a total of 91 elections = 3 ballot formats x 2 ballot aggregation methods x 3 AI models x 5 personal traits.
@@ 4.3 Evaluation of choices by AI representatives
The emulated elections with AI representatives are compared to the real-world elections of human voters at two levels: (i) individual choice, i.e. the ballots, and (ii) collective choice, i.e. the resulting voting outcomes. Consistency is the main measure of assessment.
@@ 4.4 Human cognitive biases in AI collective decision making
Surrogation biases. This reflects how humans favor simpler measures to assess the impact over ones that are more precise and harder to evaluate. Korteling et al. [41] argue that these biases manifest when deciding projects with large societal impact such as health or education, while their outcome is subject of different satisfaction levels among citizens. The project outcomes may be perceived as successful by part of society using easy-to-evaluate metrics instead of looking at long-term effects on the community. For instance, a timely vaccination drive may be preferred over significant changes to vaccination protocols or health insurance policies covering vaccination. Hence these projects are likely to be more preferred as they come with more intuitive ways to assess for the broader population. This is reflected by the average winning rate of the health and education-related projects of 38.1% and 36.2% respectively in Poland [39], where participatory campaigns have been actively hosted in the last decade.